Sol Summary – May 6th

Crew 265 Sol Summary Report 06-05-2022

Sol: 12

Summary Title: Return to Earth

Author’s Name: Marc Levesque, Crew Commander

Mission Status: Nominal

Sol Activity Summary:

Our crew woke up this morning to the crisp desert air of Earth. The first order of business was to retrieve the repeater on North Ridge, a much faster process than setting it up in EVA suits. The rest of the sol was spent packing our equipment and personal gear, finishing reports, starting to clean the station, and meeting with the Station Director to review the mapping and communications projects.

After a busy sol, we enjoyed one last dinner together.

Look ahead plan: Tomorrow we’ll finish cleaning the Hab and then depart for our respective homes or other destinations. We’re tired but satisfied with what we have accomplished.

Anomalies in work: None

Crew Physical Status: Good

EVA: None

Reports to be filed:

Sol Summary

Operations Report

Support Requested: None

Operations Report – May 6th

Crew 265 Operations Report Sol 12 06-05-2022

SOL: 12

Name of person filing report: Dave Laude/Sergii Iakymov

Non-nominal systems: Furnace. Room#2 doorknob. Percy’s left rear tail light is non-functional but the brake lights are working. Suit #6. Rover Perseverance.

Notes on non-nominal systems: Furnace is functional, but circuit breaker (labeled "air handler") is now off. The doorknob feels like it’s about to fail. Upon today’s inspection, we found Rover Perseverance has cracking plastic exterior side panels on both sides. They are located just below the seat bottoms.

Spirit rover used: No
Hours: 179.5
Beginning charge: 100
Ending charge: 100
Currently charging: Yes

Opportunity rover used: No
Hours: 97.6
Beginning charge: 100
Ending charge: 100
Currently charging: Yes

Curiosity rover used: No
Hours: 190.6
Beginning charge: 100
Ending charge: 100
Currently charging: Yes

Perseverance rover used: No
Hours: 239.6
Beginning charge: 100
Ending charge: 100
Currently charging: Yes

General notes and comments:

Summary of Hab operations: Executing end of sim cleanup checklist and Food Inventory list.

Water (static tank): Not full. ~167 gallons.
Water (loft tank): 30 gallons
Water Meter: 158357.4 units

Static to Loft Pump used – No
Static tank pipe heater (on or off): Off
Static tank heater (On or off): Off
Toilet tank emptied: No

Summary of internet: Nominal

Summary of suits and radios: No known issues. Batteries charging.

Summary of GreenHab operations: NA


Heater: N/A

Supplemental light: N/A

Harvest: None

Summary of ScienceDome operations: Not used.

Dual split: Heat or AC, On or Off

Summary of RAM operations: None

Summary of any observatory issues: NA

Summary of health and safety issues: All crew members are feeling good.

Questions, concerns and requests to Mission Support: Thank you Mission Support for your service during our rotation!

Copyright © The Mars Society. All rights reserved. | Main Site